### v2

# Efficient Long Sequential Low-rank Adaptive Attention for Click-through rate Prediction

Abstract.
In the context of burgeoning user historical behavior data, Accurate click-through rate(CTR) prediction requires effective modeling of lengthy user behavior sequences. As the volume of such data keeps swelling, the focus of research has shifted towards developing effective long-term behavior modeling methods to capture latent user interests. Nevertheless, the complexity introduced by large scale data brings about computational hurdles. There is a pressing need to strike a balance between achieving high model performance and meeting the strict response time requirements of online services. While existing retrieval-based methods (e.g., similarity filtering or attention approximation) achieve practical runtime efficiency, they inherently compromise information fidelity through aggressive sequence truncation or attention sparsification. This paper presents a novel attention mechanism. It overcomes the shortcomings of existing methods while ensuring computational efficiency. This mechanism learn compressed representation of sequence with length $L$ via low-rank projection matrices (rank $r\ll L$), reducing attention complexity from $O(L)$ to $O(r)$. It also integrates a uniquely designed loss function to preserve nonlinearity of attention. In the inference stage, the mechanism adopts matrix absorption and pre-storage strategies. These strategies enable it to effectively satisfy online constraints. Comprehensive offline and online experiments demonstrate that the proposed method outperforms current state-of-the-art solutions.
Click-Through Rate Prediction; Long-term User Behavior Modeling; Low-rank Compression
copyright: none
1. Introduction
Click-Through Rate (CTR) prediction is a critical element in online advertising and recommendation systems. With the rapid growth of user behavior data, effectively modeling long-term user behavior and capturing user interests has become a focal point of research (Covington et al., 2016; Zhou et al., 2018, 2019; Pi et al., 2019). Long-term user behavior modeling provides valuable insights into user preferences and trends over extended periods (Pi et al., 2020; Chen et al., 2021; Cao et al., 2022; Chang et al., 2023; Si et al., 2024), though it simultaneously introduces challenges related to computational complexity. These challenges necessitate careful consideration of the trade-offs between the model performance and the request time limit for online services.
In addressing the issue of long sequences, the industry commonly adopts information filtering techniques to select top-k items, often based on item attributes like SIM(Pi et al., 2020), or some form of low-precision attention like ETA(Chen et al., 2021). and SDIM(Cao et al., 2022) This approach has been validated in various studies (Chang et al., 2023; Si et al., 2024; Chen et al., 2021; Cao et al., 2022; Pi et al., 2020), including UBR4CTR (Qin et al., 2020), which employs end-to-end retrieval-augmented transformers, as well as TWIN (Chang et al., 2023) and TWIN V2 (Si et al., 2024), which utilize GPU-optimized hierarchical attention to process sequences and significantly enhance CTR prediction accuracy effectively. However, this approach sacrifices some products with low correlation in some aspects. It also causes information loss. For example, when the target item is pants, the system retrieves the user’s past behaviors related to buying pants generating the top K relevant features. However, it may overlook the user’s purchases of other light-colored clothes, which implies an interest in dark-colored pants.
To overcome these limitations while ensuring efficient online computation, we explore alternative methodologies that leverage all available information. Drawing inspiration from the Multi-Head Latent Attention (MLA) (DeepSeek-AI et al., 2024) technique proposed by DeepSeek, we note that MLA efficiently compresses and decompresses features during execution, achieving better performance. It focuses on reducing the storage of the kv cache. However, it doesn’t effectively reduce computation during online inference, making it unsuitable for our requirements.
To address these challenges, we propose Low-Rank Efficient Attention(LREA), a novel attention mechanism designed specifically for modeling long-term user behavior with long-term sequences. LREA employs low-rank matrix decomposition at the sequence level, optimizing runtime performance while preserving the integrity of long-term behavior data. Furthermore, LREA integrates tailored designed loss function that maintain attention capabilities and simplify the computation. In the inference phase, it utilizes matrix fusion and pre-storage techniques to accommodate runtime constraints effectively.
In summary, We propose LREA, the first approach focused on modeling entire long-term user behavior. LREA employs low-rank matrix decomposition and matrix absorption methods to meet online inference latency
2. Preliminaries
2.1. Task Formulation
CTR prediction is formulated as a binary classification problem, wherein the objective is to learn a function $f:\mathbb{R}^{D}\rightarrow\mathbb{R}$ based on a given training dataset $\mathcal{D}=\{(x_{1},y_{1}),\ldots,(x_{|\mathcal{D}|},y_{|\mathcal{D}|})\}$. Each training sample $x_{k}\in\mathbb{R}^{D}$ is represented as a feature vector encapsulating various user, item, and contextual features.
Given the high dimensionality and sparsity of the input feature space, an embedding layer is employed. Each sparse feature is converted into a lower-dimensional dense representation through an embedding matrix. The embeddings generated from the different input features are concatenated as $x_{k}$ and input to a Multi-Layer Perceptron (MLP) layer, formalized as:
(1) $\hat{y}_{k}=\sigma(MLP(x_{k}))$
where $\sigma(\cdot)$ is the sigmoid function, and $MLP$ is the mapping function that transforms the input feature representation $x_{k}\in\mathbb{R}^{d_{0}}$ into a probability score.
The optimization objective is the binary cross-entropy loss:
(2) $\mathcal{L}=-\frac{1}{|\mathcal{D}|}\sum_{k=1}^{|\mathcal{D}|}\left(y_{k}\log(\hat{y}_{k})+(1-y_{k})\log(1-\hat{y}_{k})\right),$
where $\mathcal{D}$ represents the training dataset, and $y_{k}\in\{0,1\}$ indicates the ground truth label for the corresponding interaction.
2.2. long-term User Behavior Sequence
Behavior sequences provide a comprehensive view of a user’s historical interactions with various items. For each user $u\in U$, we can model their behavioral history as a sequence of actions $\mathcal{H}_{su}=[s_{1},s_{2},\ldots,s_{L}]$, where $s_{i}$ denotes the $i$-th interacted item. After feeding into embedding layer,the $i$-th interacted item’s representation is $e_{i}\in\mathbb{R}^{1\times d}$ where d is dimension of a item’s representation. the Multi-head attention is widely applied on modeling user behavior sequence (Chen et al., 2019; Xiao et al., 2017; Xu et al., 2020; Zhang et al., 2024). In CTR prediction task, target item acts as query ($\boldsymbol{Q}$) and each item in user behavior sequence acts as key ($\boldsymbol{K}$) and value ($\boldsymbol{V}$). We call this multi-head attention structure as multi-head target attention, which is abbreviated as TA. Let $E_{t}\in\mathbb{R}^{1\times d}$, $E_{s}\in\mathbb{R}^{L\times d}$ represent input embedding matrices for target item and behavior sequence. we have $E_{s}=[e_{1},e_{2},\ldots,e_{L}]^{T}$.
The online inference complexity of multi-head target attention is $O(L*B*d)$, where $B$ is the number of online inference candidate items and $L$ is the length of the long-term user behavior sequence. As the length of the behavior sequence increases from tens to thousands, the Online computation latency is not tolerable.
3. Low-Rank Efficient Attention
Figure 1. The structure of LREA. In the off-line training stage, LREA updates matrices $W_{Comp}$ and $W_{Decomp}$ without computational optimization, subsequently caching them with $E_{s}$ and $E_{s}^{T}$ in HBM. During offline inference, LREA utilizes compressed sequences $E_{Auxabsorb}$ and $E_{Comp}$, which reduce the length of user behavior sequence, enhancing the computation efficiency.
[…]
Diagram illustrating the LREA structure, showing training and inference stages with compressed sequences and matrices for efficient computation.
In this section, we introduce LREA, a novel attention mechanism designed to tackle the inference runtime limitations of long-term user behavior sequences in CTR prediction tasks. Previous works mainly focused on top K retrieval sequences, reducing their length from $\mathbb{R}^{L}$ to $\mathbb{R}^{K}$. Recent low-rank matrix decomposition techniques, such as LoRA and MLA, have shown effectiveness in related fields. LREA builds on this by applying low-rank matrix decomposition to user behavior sequences and using matrix absorption techniques to address online runtime limitations.
3.1. Long Sequence Compression
CTR prediction fundamentally relies on the interaction of user behavior sequences and target items. Let $\mathcal{H}_{lu}=[s_{1},s_{2},\ldots,s_{L}]$ denote a user’s long-term behavior sequence, and $E_{t}\in\mathbb{R}^{d}$ represent the embedding of the target item:
(3) $TA(E_{s},E_{t})=E_{s}^{T}\cdot AttScore(E_{s},E_{t})$
In this equation, $E_{s}^{T}$ is the transposed matrix of $E_{s}$, and $AttScore(E_{s},$ $E_{t})\in\mathbb{R}^{L\times 1}$ represents the attention score designed to the behavior sequence of length $L$, which is a critical step in CTR prediction. However, the calculation of target attention becomes a bottleneck during inference due to the constraints posed by the lengths of long-term sequences.
Low-rank Matrix Decompression To mitigate information loss associated with pure compression, our approach, LREA, employs principles of low-rank matrix decomposition(Hu et al., 2022). Specifically, it utilizes two matrices $W_{comp}$ and $W_{Decomp}$ to compress and subsequently decompress the original sequence representation, as follows:
(4) $E_{s}^{LREA}=(E_{s}^{T}W_{Comp}W_{Decomp})^{T}$
Here, $W_{Comp}\in\mathbb{R}^{L\times r}$ and $W_{Decopm}\in\mathbb{R}^{r\times L}$, where $L$ is the original sequence length and $r$ is the reduced rank. $E_{comp}=E_{s}^{T}W_{Comp}\in\mathbb{R}^{d\times r}$ can be seen as a compressed representation of the behavior sequence. This operation helps to preserve essential information while handling the compressive aspect of the sequence. We can express target attention computation like implementation of DIN target attention(Zhou et al., 2018) as:
(5) $\begin{split}&~{}~{}~{}~{}~{}~{}\phi\left[\bigg{(}E_{t}\oplus W_{Decomp}^{T}E_{Comp}^{T}\oplus W_{Decomp}^{T}(E_{t}\odot E_{Comp}^{T})\bigg{)}W_{in}\right]W_{o}\\ &=[\phi(E_{t}W_{1})\ +\phi(W_{Decomp}^{T}E_{Comp}^{T}W_{2})+\phi(W_{Decomp}^{T}(E_{t}\odot E_{Comp}^{T})W_{3})]W_{o}\end{split}$
where $W_{in},W_{o}$ are linear matrices and $W_{in}$ could split into three parts $W_{1},W_{2},W_{3}$ to calculate separately. $\oplus$ means vector concatenation operation, $\odot$ means hadamard production, $\phi(\cdot)$ is a nonlinear activation function. Here we will emphasize the $(E_{s}\odot E_{t})$ formulation’s signification for interaction of the target item, and sequence items’ representation and nonlinearity introduced by $\phi(\cdot)$ activation function. These two key components have the same function correspond to the inner product $QK^{T}$ and softmax in multi-head target attention respectively.
At this point, We want to split $E_{comp}={\bigg{(}E_{s}^{T}W_{Comp}\bigg{)}}_{(d\times r)}$ out because it compression the $L$-dimension of $E_{s}$ to a low-rank $r$ the and at the same time to keep the information integrity of the formula above.
Matrix Absorption First we take the key part $\phi(W_{Decomp}^{T}E_{Comp}^{T}\times W_{2})$ alone to analyze, if we can find a $\phi(\cdot)$ satisfy:
(6) $\phi(AB)\ =A\phi(B)$
Then we could split out $E_{comp}={\bigg{(}E_{s}^{T}W_{Comp}\bigg{)}}_{(d\times r)}$ for cache to simplify the computation for target attention. A strict solution of this equation is $\phi(x)=kx,k\in\mathbb{R}$. But this will cause the network to lose nonlinearity. A potential solution of nonlinear activation function $\phi(\cdot)$ to approximate $\phi(x)=kx$ is function of the ReLU function:
(7) $ReLU(x)=\begin{cases}x&\text{if }x\geq 0\\ 0&\text{if }x<0\end{cases}$
Because ReLU mostly meets the requirements of the solution to $\phi(AB)\ =A\phi(B)$, and is more suitable than other activation functions. We could choose from the function of ReLU series. Leaky ReLU and Parametric ReLU(He et al., 2015) will be $ax$ if $x<0$. Here we use Leaky ReLU. No matter what variant ReLU function we need to guarantee to the greatest extent that the elements of matrix $A$ and $B$ need to have consistent signs, or both be non-negative. But if both $A$ and $B$ are non-negative, The formula is not nonlinear anymore, this will lead to sub-optimal performance for attention mechainism. So there is a balance here between non-negativity and non-linearity.
This allows us to derive:
(8) $\begin{split}\phi(W_{Decomp}^{T}E_{Comp}^{T}W_{2})&=W_{Decomp}^{T}\times\phi(E_{Comp}^{T}W_{2})\\ \phi(W_{Decomp}^{T}(E_{t}\odot E_{Comp}^{T})W_{3})&=W_{Decomp}^{T}\times\phi((E_{t}\odot E_{Comp}^{T})W_{3})\end{split}$
Thereby, the original target attention computation to our proposed low-rank efficient attention could be approximated as:
(9) $\begin{split}LREA(E_{s}^{T},E_{t})=E_{S}^{T}W_{Decomp}^{T}\phi\bigg{(}[E_{t}\oplus W_{comp}^{T}\oplus(E_{t}\odot W_{comp}^{T})]W_{in}\bigg{)}W_{o}\end{split}$
The $W_{Decomp}$ could be absorbed by $E_{auxabsorb}=E_{S}^{T}W_{Decomp}^{T}\in\in\mathbb{R}^{d\times r}$, just like $E_{comp}$, It compression the L-dimension of $E_{s}$ to a low-rank $r$ Finally the output of LREA will be concatenated with all other features’ embedding and then fed into MLP to predict $\hat{y}$.
3.2. Model Training and Inference
To ensure the validity of Eq. 6, we need to impose non-negativity constraints on matrices $A$ and $B$. Therefore, we introduce a non-negativity loss function:
(10) $\mathcal{L}_{non-neg}=||max(0,-W_{Decomp}^{T})||_{2}^{2}\ +\ ||max(0,-W_{comp}^{T}E_{s})||_{2}^{2}$
where $||A||_{2}=\sqrt{\sum_{i,j}|A_{ij}}|^{2}$ is the Frobenius norm of a matrix, so $\mathcal{L}_{non-neg}$ will only impose a penalty on the negative elements in the matrix, while the penalty term for positive elements is 0. The final training objective function is derived by
(11) $\mathcal{L}_{LREA}=\mathcal{L}\ +\lambda\ \mathcal{L}_{non-neg}$
Table 1. Time complexity of different methods at training and online-serving stage.
Method Training Online-Serving
DIN (Zhou et al., 2018) O(BLd) O(BLd)
SIM (Pi et al., 2020) O(Blog(M) + Bkd) O(Blog(M) + Bkd)
ETA (Chen et al., 2021) O(Lmlog(d) + BLm + Bkd) O(BLm + Bkd)
SDIM (Cao et al., 2022) O(Lmlog(d) + Bmlog(d)) O(Bmlog(d))
LREA O(BLrd) O(Brd)
where $\lambda$ is the hyper-parameter to balance the nonlinearity and non-negativity.
Fig. 1 illustrates the overview of LREA. During the offline training stage, LREA trains and updates two additional matrices, $W_{Compress}$ and $W_{Decompress}$, without performing any optimization on computation. Once the training phase is complete, these two matrices are multiplied with the $E_{S}$ and the $E_{S}^{T}$, respectively, cached in the High Bandwidth Memory (HBM) denoted as $E_{Compress}$ and $E_{Auxabsorb}$. In the online inference stage, LREA does not leverage the original long-term user behavior sequences. Instead, it utilizes the pre-stored sequences $E_{Auxabsorb}$ and $E_{Comp}$, in which the sequence length has been effectively reduced through absorption techniques. This method accelerates the inference process, ensuring a more efficient deployment of the model in a real-world system. By focusing on the compressed representations, LREA can deliver faster results while maintaining the integrity of the user behavior data, ultimately enhancing the overall performance of the CTR model.
Table 1 shows the computing complexity of LREA compared with prior works. In this analysis, the relevant parameters for our analysis include $B$, the number of candidate items for each request; $m$, the number of hashes employed; $L$, $k$ and $r$, representing the lengths of the original, retrieved user behavior sequences and compressed user behavior sequences, respectively; $M$, the size of the attribute inverted index in the SIM; and $d$, the model’s hidden size.
4. Experiments
In this section, we describe the experimental setup and conduct extensive experiments on both public and industrial dataset to evaluate the performance and the cost of the proposed LERA model.
4.1. Experiments Setup
Datasets. Our offline model evaluation are conducted on two widely used CTR prediction datasets for research and one real-world industrial dataset.
Amazon dataset [^1] https://cseweb.ucsd.edu/~jmcauley/datasets.html#amazon_reviews. It is derived from user interactions on Amazon’s e-commerce platform(Ni et al., 2019). We use Books collection of reviews, pre-process the data follow (Pi et al., 2020). the recent 10 user behaviors as short-term user sequential features and recent 90 user behaviors as long-term user sequential features.
Taobao dataset [^2] https://tianchi.aliyun.com/dataset/dataDetail?dataId=649. It is based on user behavior logs from Taobao, one of the major e-commerce platform in China and contains 89 million records. we take the click behaviors for each user and sort them according to time in an attempt to construct the behavior sequence.The we use the first 64 behaviors as short-term behavior sequence and the rest of 500 clicked items as long-term sequence features.
Industrial dataset. It is collected from the logs of our real-world industrial e-commerce advertising platform. The dataset encompasses advertising impression logs and corresponding click labels during the period from December 3, 2024, to December 23, 2024.It cotains 2 billion records from 10 million users. The user historical clicked items reach tens of thousands within the past year. We use the recent 64 click behaviors as short-term sequential features and 10000 items as long-term sequential features.
Table 2. Results on public datasets and industrial dataset
Models Amazon Taobao Industrial dataset
$AUC$ $AUC$ $AUC$ $GAUC$
DIN(Long sequence) 0.8101 0.8903 0.7149 0.6337
DIN(W/o long seq.) 0.7833 0.8535 0.7005 0.6187
SIM 0.7919 0.8667 0.7031 0.6208
ETA 0.7984 0.8651 0.7083 0.6235
SDIM 0.8013 0.8734 0.7097 0.6239
TWIN 0.8015 0.8838 0.7103 0.6258
LREA 0.8098 0.8904 0.7154 0.6343
Competitors. We compare the proposed LERA method with mainstream CTR models and various of two-stage models for long-term user behaviors. DIN(W/o long seq.) (Zhou et al., 2018): Only short-term sequences are involved. DIN(Long sequence): We apply the DIN target attention for both short-term and long-term sequences, but this method is not possible to deploy online for the intolerable inference latency. We also implement baseline of two-stage method for long-term user behaviors: SIM (Pi et al., 2020), ETA (Chen et al., 2021), SDIM (Cao et al., 2022), TWIN (Chang et al., 2023).
Parameter Setting. The hidden layers’ sizes are set to a fixed tuple [512, 256, 128]. Each hidden layer follows a Leaky ReLU (Maas et al., 2013) and a BatchNorm Layer (Ioffe, 2015) to accelerate the model training. All methods are trained with Adagrad for optimization. We tuned the learning rate in 0.01, 0.005, 0.001. The batch size is set to 1024 and embedding dimension $d$ is set to 16.
Evaluate Metrics. we use two widely adopted metrics: AUC and GAUC. GAUC performs a weighted average over all users’ request group, and the weights are set to the number of samples of this user’s request..We will only apply GAUC on industrial dataset due to the number of groups is relatively small in public datasets.
Table 3. Hyper-parameters Analysis on Industrial Dataset
$\lambda=0.3$ $r=128$
LREA AUC GAUC LREA AUC GAUC
$r=32$ 0.7081 0.6222 $\lambda=0.05$ 0.7121 06298
$r=64$ 0.7097 0.6254 $\lambda=0.1$ 0.7138 06312
$r=128$ 0.7154 0.6343 $\lambda=0.3$ 0.7154 0.6343
- - $\lambda=0.5$ +0.7137 0.6315
- - $\lambda=0.8$ +0.7129 0.6297
4.2. Experiment Performance
Overall Perfomance. The evaluation results is shown in Table 2. SIM, EAT, SDIM, TWIN, outperform DIN. That indicates long-term behavior sequence modeling is helpful for more accurate modeling of CTR since it can cover more user’s historical interests. DIN(Long Sequence) has a stable improvement of AUC gain of 0.86%/0.65% compared with the best two-stage method on Amazon and Taobao dataset. We can also see a AUC gain of 0.49% and a GAUC gain of 0.79% in industrial dataset when comparing DIN(Long Sequence) with TWIN. But due to the limitation of online inference latency, we can’t deploy DIN to long sequence for modeling long-term user interests. However, our LREA also outperforms all two-stage methods and it achieves similar results to long sequence target attention, and even exceeds DIN on industrial datasets but its online computation complexity is much less than that of long sequence DIN.
Online A/B Test. We conduct online A/B experiments to evaluate LREA in our online advertising system. The experiment lasts for 12 days and LREA achieves 5.88% CTR and 4.26% RPM gain compared to product model (DIN+SIM).
Hyper-parameters Analysis. We conduct analysis experiments of two important hyper-parameters on our industrial dataset: compression rank $r$ and non-negative loss coefficient $\lambda$. As shown in Table 3. First we fix $\lambda$ to 0.3, tune r in from 32 to 128. The Compression rank of $r=128$ achieves the best performance, and AUC/GAUC drop significantly when $r$ becomes smaller. That indicates that we can’t compression the representation of the original long sequences to a very low rank. We will not test $r$ 128, because although the larger r is, the better performance we get, there should be a appropriate rank for compression to balance cache memory size and model performance since we can’t deploy it online due to model inference latency.
As shown in the second major column of Table 3, we fix $r$ at 128 and adjust the value of $\lambda$ between 0.05 and 0.8. As mentioned earlier, the $\lambda$ parameter controls the nonlinearity and non-negativity in the compression module. If $\lambda$ is too large, it reduces the nonlinearity, resulting in a decline in model performance. Conversely, if $\lambda$ is too small, it skews the loss of the main task of CTR estimation. In our experiments, the model performs best when $\lambda=0.3$.
5. Conclusion
In this work, we introduced LREA, a novel attention mechanism that enhances CTR prediction in online advertising and recommendation systems. LREA effectively integrates long-term user behavior by using low-rank matrix decomposition and matrix absorption. This mitigates the performance degradation caused by traditional top K selection. Experimental results demonstrate that our model achieves significant improvements over the competitors.