### v4

# LREA: Low-Rank Efficient Attention on Modeling Long-Term User Behaviors for CTR Prediction

Abstract.
With the rapid growth of user historical behavior data, user interest modeling has become a prominent aspect in Click-Through Rate (CTR) prediction, focusing on learning user intent representations. However, this complexity poses computational challenges, requiring a balance between model performance and acceptable response times for online services. Traditional methods often utilize filtering techniques. These techniques can lead to the loss of significant information by prioritizing top K items based on item attributes or employing low-precision attention mechanisms. In this study, we introduce LREA, a novel attention mechanism that overcomes the limitations of existing approaches while ensuring computational efficiency. LREA leverages low-rank matrix decomposition to optimize runtime performance and incorporates a specially designed loss function to maintain attention capabilities while preserving information integrity. During the inference phase, matrix absorption and pre-storage strategies are employed to effectively meet runtime constraints. The results of extensive offline and online experiments demonstrate that our method outperforms state-of-the-art approaches.
Click-Through Rate Prediction; Long-term User Behavior Modeling; Low-rank Compression
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italy.booktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italyisbn: 979-8-4007-1592-1/25/07doi: 10.1145/3726302.3730228ccs: Information systems Recommender systems
1. Introduction
Click-Through Rate (CTR) prediction is pivotal in online advertising and recommendation systems. Previous studies on short sequences have mainly focused on the interaction between sequences and targets or trigger item (Zhou et al., 2018, 2019; Lou et al., 2024; Shen et al., 2022). With the rapid growth of user behavior data, effectively modeling long-term user behavior has become a research focus (Covington et al., 2016; Pi et al., 2019). While long-term behavior modeling offers insights into user preferences (Pi et al., 2020; Chen et al., 2021; Cao et al., 2022; Chang et al., 2023; Si et al., 2024). it also brings computational complexity, requiring a balance between model performance and online-service response time.
The industry typically adopts information filtering techniques to select the top K items, based on item attributes such as SIM (Pi et al., 2020), or applying low-precision attention (e.g., ETA (Chen et al., 2021) and SDIM (Cao et al., 2022)) to handle long sequences. This approach has been validated in various studies (Qin et al., 2020), including TWIN (Chang et al., 2023) and TWIN v2 (Si et al., 2024), which utilize GPU-optimized hierarchical attention to process sequences effectively, significantly enhancing CTR prediction accuracy. However, this approach sacrifices some items with low correlation. For example, when the target item is pants, the system focuses on pants-related behaviors for top K features but may overlook light-colored clothing purchases, which may imply an interest in dark pants.
To overcome these limitations and ensure efficient online computation, we explore alternative methodologies leveraging all data. Inspired by DeepSeek’s MLA (DeepSeek-AI et al., 2024), which compresses hidden representations for better performance and reduces KV cache storage. However, We find it unsuitable for our requirements as it does not reduce inference computation.
We propose Low-Rank Efficient Attention(LREA), a novel attention mechanism specifically designed for modeling long-term user behavior. LREA employs low-rank matrix decomposition at the sequence length level, optimizing performance while preserving the integrity of long-term behavior data. Furthermore, LREA integrates a tailor-designed loss function that maintains attention capabilities and simplifies the computation. In the inference phase, it utilizes matrix absorption and pre-storage techniques to effectively accommodate runtime constraints.
In summary, we propose LREA, the first approach focused on modeling entire long-term user behavior. LREA employs low-rank matrix decomposition and matrix absorption methods to meet online inference latency.
2. Preliminaries
2.1. Task Formulation
CTR prediction is formulated as a binary classification problem, wherein the objective is to learn a function $f:\mathbb{R}^{D}\rightarrow\mathbb{R}$ based on a given training dataset $\mathcal{D}=\{(x_{1},y_{1}),\ldots,(x_{|\mathcal{D}|},y_{|\mathcal{D}|})\}$. Each sparse feature of a training sample is converted into a low-dimensional dense representation through an embedding matrix. The embeddings generated from different features are concatenated as $x_{k}\in\mathbb{R}^{D}$ and fed to a Multi-Layer Perceptron (MLP) layer, formalized as:
(1) $\hat{y}_{k}=\sigma(MLP(x_{k}))$
where $\sigma(\cdot)$ is the sigmoid function, and $MLP$ is the mapping function that transforms the input $x_{k}$ into a probability score. The optimization objective is the cross-entropy loss:
(2) $\mathcal{L}=-\frac{1}{|\mathcal{D}|}\sum_{k=1}^{|\mathcal{D}|}\left(y_{k}\log(\hat{y}_{k})+(1-y_{k})\log(1-\hat{y}_{k})\right)$
where $\mathcal{D}$ represents the training dataset, and $y_{k}\in\{0,1\}$ indicates the ground truth label for the corresponding interaction.
2.2. long-term User Behavior Sequence
Behavior sequences provide a comprehensive view of users’ historical interactions with various items. For each user $u\in U$, we can model their behavioral history as a sequence of actions $\mathcal{H}_{lu}=[s_{1},s_{2},\ldots,s_{L}]$, where $s_{i}$ denotes the $i$-th interacted item. After being fed into the embedding layer, the representation of the $i$-th interacted item is $e_{i}\in\mathbb{R}^{1\times d}$ where d is the dimension of the representation of an item. Multi-head attention is widely applied to model user behavior sequence (Chen et al., 2019; Xiao et al., 2017; Xu et al., 2020; Zhang et al., 2024). In CTR prediction tasks, target item acts as query ($\boldsymbol{Q}$) and each item in user behavior sequence acts as key ($\boldsymbol{K}$) and value ($\boldsymbol{V}$). We call this multi-head attention structure as multi-head target attention, which is abbreviated as MHTA. The online inference complexity of multi-head target attention is $O(L*B*d)$, where $B$ is the number of online inference candidate items and $L$ is the length of the long-term user behavior sequence. As the length of the behavior sequence increases from tens to thousands, the Online computation latency is not tolerable.
3. Low-Rank Efficient Attention
Figure 1. The structure of LREA. In the off-line training stage, LREA updates matrices $W_{Comp}$ and $W_{Decomp}$ without computational optimization, subsequently caching them with $E_{s}$ and $E_{s}^{T}$ in HBM. During offline inference, LREA utilizes compressed sequences $E_{Auxabsorb}$ and $E_{Comp}$, which reduce the length of user behavior sequence, enhancing the computation efficiency.
[…]
Diagram illustrating the LREA structure, showing training and inference stages with compressed sequences and matrices for efficient computation.
In this section, we introduce LREA, a novel attention mechanism designed to tackle the inference limitations of modeling long-term user behavior sequences in CTR prediction. LREA applies low-rank matrix decomposition to user behavior sequences and uses matrix absorption techniques to address online runtime limitations.
3.1. Long Sequence Compression
CTR prediction fundamentally relies on the interaction of user behavior sequences and target items. Let $\mathcal{H}_{lu}=[s_{1},s_{2},\ldots,s_{L}]$ denote a user’s long-term behavior sequence, and $E_{t}\in\mathbb{R}^{d}$ represent the embedding of the target item:
(3) $TA(E_{s},E_{t})=E_{s}^{T}\cdot AttScore(E_{s},E_{t})$
In this equation, $E_{s}=[e_{1},e_{2},\ldots,e_{L}]^{T}$ is the embedding representation of behavior sequence and $AttScore(E_{s},$ $E_{t})\in\mathbb{R}^{L\times 1}$ represents the attention score designed to the behavior sequence.
Low-rank Matrix Decompression To mitigate information loss associated with pure compression, our approach, LREA, employs principles of low-rank matrix decomposition (Hu et al., 2022). Specifically, it utilizes two matrices $W_{Comp}$ and $W_{Decomp}$ to compress and subsequently decompress the original sequence representation:
(4) $E_{s}^{LREA}=(E_{s}^{T}W_{Comp}W_{Decomp})^{T}$
Here, $W_{Comp}\in\mathbb{R}^{L\times r}$ and $W_{Decomp}\in\mathbb{R}^{r\times L}$, where $L$ is the original sequence length and $r$ is the reduced rank. $E_{Comp}=E_{s}^{T}W_{Comp}\in\mathbb{R}^{d\times r}$ can be seen as a compressed representation of the behavior sequence. This operation helps to preserve essential information while handling the compressive aspect of the sequence. First we can formulate attention score computation of DIN (Zhou et al., 2018) as:
(5) $\begin{split}AttScore_{DIN}=&\phi\left[\bigg{(}E_{t}\oplus E_{s}\oplus(E_{t}\odot E_{s})\bigg{)}W_{in}\right]W_{o}\\ =&\phi[(E_{t}W_{1})+(E_{s}W_{2})+(E_{t}\odot E_{s})W_{3}]W_{o}\end{split}$
where $W_{in},W_{o}$ are linear matrices and $W_{in}$ could split into three parts $W_{1},W_{2},W_{3}$ to calculate separately. $\oplus$ means vector concatenation operation, $\odot$ means hadamard production, $\phi(\cdot)$ is a nonlinear activation function. Here we will emphasize the $(E_{t}\odot E_{s})$ formulation’s signification for interaction of the target item and sequence items’ representation, and nonlinearity introduced by $\phi(\cdot)$. These two components have the same function corresponding to the inner product $QK^{T}$ and softmax in MHTA respectively. Here we could introduce interaction more simply by $(E_{t}\odot E_{Comp}^{T})$, Then if we replace $E_{s}$ in DIN attention with $E_{s}^{LREA}$ in Eq. 4, replace $(E_{s}\odot E_{t})$ with $W_{Decomp}^{T}(E_{t}\odot E_{Comp}^{T})$, at the same time replace $E_{t}$ as $W_{Decomp}^{T}E_{t}$(after broadcast $E_{t}\in\mathbb{R}^{r\times d}$), we transform DIN attention score in formula 5 into LREA score $AttScore_{LREA}$:
(6) $\begin{split}&AttScore_{LREA}\\ =&\phi[W_{Decomp}^{T}E_{t}W_{1}\ +W_{Decomp}^{T}E_{Comp}^{T}W_{2}+W_{Decomp}^{T}(E_{t}\odot E_{Comp}^{T})W_{3}]W_{o}\\ =&\phi[W_{Decomp}^{T}(E_{t}W_{1}\ +E_{Comp}^{T}W_{2}+E_{t}\odot E_{Comp}^{T}W_{3})]W_{o}\\ \end{split}$
Matrix Absorption If we can find a $\phi(\cdot)$ satisfy:
(7) $\phi(AB)\ =A\phi(B)$
we can easily derive this formula:
(8) $\begin{split}LREA(E_{s}^{T},E_{t})&=E_{s}^{T}\cdot AttScore_{LREA}\\ &=E_{S}^{T}W_{Decomp}^{T}\phi\bigg{(}[E_{t}\oplus E_{Comp}^{T}\oplus(E_{t}\odot E_{Comp}^{T})]W_{in}\bigg{)}W_{o}\end{split}$
The $W_{Decomp}$ could be absorbed by $E_{auxabsorb}=\bigg{(}E_{S}^{T}W_{Decomp}^{T}\bigg{)}_{(d\times r)}$. just like $E_{Comp}$, It can compress the L-dimension of $E_{s}$ to a low-rank $r$. Finally, the output of LREA will be concatenated with all other features’ embedding and then fed into MLP to predict $\hat{y}$.
A strict solution of $\phi(AB)\ =A\phi(B)$ is $\phi(x)=kx,k\in\mathbb{R}$. But this will cause the network to lose nonlinearity. A potential solution of nonlinear activation function $\phi(\cdot)$ to approximate $\phi(x)=kx$ is the ReLU function:
(9) $ReLU(x)=\begin{cases}x&\text{if }x\geq 0\\ 0&\text{if }x<0\end{cases}$
Because ReLU mostly meets the requirements of the solution to $\phi(AB)\ =A\phi(B)$, and is more suitable than other activation functions. We could choose from the functions of ReLU series, Leaky ReLU, and Parametric ReLU(He et al., 2015). Here we use Leaky ReLU. No matter what variant ReLU function we choose, we need to guarantee to the greatest extent that the elements of matrix $A$ and $B$ have consistent signs, or both are non-negative. But if both $A$ and $B$ are non-negative, the formula is not nonlinear anymore. This will lead to sub-optimal performance for the attention mechanism. So there is a balance here between non-negativity and non-linearity.
3.2. Model Training and Inference
To ensure the validity of Eq. 7, we need to impose non-negativity constraints on matrices $A$ and $B$. Therefore, we introduce a non-negativity loss function:
(10) $\mathcal{L}_{non-neg}=||max(0,-W_{Decomp}^{T})||_{2}^{2}\ +\ ||max(0,-W_{Comp}^{T}E_{s})||_{2}^{2}$
where $||A||_{2}=\sqrt{\sum_{i,j}|A_{ij}}|^{2}$ is the Frobenius norm of matrix A, so $\mathcal{L}_{non-neg}$ will only impose a penalty on the negative elements in the matrix, while the penalty term for positive elements is 0. The final training objective function is derived by:
(11) $\mathcal{L}_{LREA}=\mathcal{L}\ +\lambda\ \mathcal{L}_{non-neg}$
Table 1. Time complexity of different methods at training and online-serving stage.
Method Training Online-Serving
DIN (Zhou et al., 2018) O(BLd) O(BLd)
SIM (Pi et al., 2020) O(Blog(M) + Bkd) O(Blog(M) + Bkd)
ETA (Chen et al., 2021) O(Lmlog(d) + BLm + Bkd) O(BLm + Bkd)
SDIM (Cao et al., 2022) O(Lmlog(d) + Bmlog(d)) O(Bmlog(d))
LREA O(BLrd) O(Brd)
Where $\lambda$ is a hyper-parameter to balance the nonlinearity and non-negativity.
Fig. 1 illustrates the overview of LREA. During the offline training stage, LREA trains and updates two additional matrices, $W_{Compress}$ and $W_{Decompress}$, without performing any optimization on computation. Once the training phase is complete, these two matrices are multiplied with $E_{S}$ and $E_{S}^{T}$, respectively, cached in the High Bandwidth Memory (HBM) denoted as $E_{Compress}$ and $E_{Auxabsorb}$. In the online inference stage, LREA does not leverage the original long-term user behavior sequences. Instead, it utilizes the pre-stored sequences $E_{Auxabsorb}$ and $E_{Comp}$, in which the sequence length has been effectively reduced through absorption techniques. This method accelerates the inference process, ensuring a more efficient model deployment in a real-world system. By focusing on the compressed representations, LREA can deliver faster results while maintaining the integrity of the user behavior data, ultimately enhancing the overall performance of the CTR model.
Table 1 shows the compute complexity of LREA compared with prior works. In this analysis, the relevant parameters for our analysis include $B$, the number of candidate items for each request; $m$, the number of hashes employed; $L$, $k$ and $r$, representing the lengths of the original, retrieved user behavior sequences and compressed user behavior sequences, respectively; $M$, the size of the attribute inverted index in the SIM; and $d$, the model’s hidden size.
4. Experiments
In this section, we describe the experimental setup and conduct extensive experiments on public and industrial datasets to evaluate the performance and cost of the proposed LERA model.
4.1. Experiments Setup
Datasets. Our offline model evaluations are conducted on two widely used CTR prediction datasets and one industrial dataset.
Amazon dataset [^1] https://cseweb.ucsd.edu/~jmcauley/datasets.html#amazon_reviews. It is derived from user interactions on Amazon’s e-commerce platform(Ni et al., 2019). We use Books collection of reviews, pre-process the data follows (Pi et al., 2020). The recent 10 user behaviors are short-term user sequential features and recent 90 user behaviors are long-term user sequential features.
Taobao dataset [^2] https://tianchi.aliyun.com/dataset/dataDetail?dataId=649. It is based on user behavior logs from Taobao, one of the major e-commerce platform in China and contains 89 million records. We take the click behaviors for each user and sort them according to time to construct the behavior sequence. Then we use the first 64 behaviors as short-term behavior sequential and the rest of 500 clicked items as long-term sequential features.
Industrial dataset. It is collected from the logs of our real-world industrial e-commerce advertising platform. The dataset encompasses advertising impression logs and the corresponding click labels during the period from December 3, 2024, to December 23, 2024. It contains 2 billion records from 10 million users. The user historical clicked items reach tens of thousands within the past year. We use the recent 64 click behaviors as short-term sequential features and 10000 items as long-term sequential features.
Table 2. Results on public datasets and industrial dataset
Models Amazon Taobao Industrial dataset
$AUC$ $AUC$ $AUC$ $GAUC$
DIN(Long sequence) 0.8101 0.8903 0.7149 0.6337
DIN(W/o long seq.) 0.7833 0.8535 0.7005 0.6187
SIM 0.7919 0.8667 0.7031 0.6208
ETA 0.7984 0.8651 0.7083 0.6235
SDIM 0.8013 0.8734 0.7097 0.6239
TWIN 0.8015 0.8838 0.7103 0.6258
LREA 0.8098 0.8904 0.7154 0.6343
Competitors. We compare LERA with the mainstream CTR models as follows: DIN(W/o long seq.) (Zhou et al., 2018): Only short-term sequences are involved. DIN(Long sequence): We apply the DIN target attention for both short and long-term sequences, but this method is not possible to deploy online due to intolerable inference latency. We also implement a two-stage method baseline for long-term user behaviors: SIM (Pi et al., 2020), ETA (Chen et al., 2021), SDIM (Cao et al., 2022), TWIN (Chang et al., 2023).
Parameter Setting. The hidden layer sizes are set to a fixed tuple [512, 256, 128]. Each hidden layer follows a Leaky ReLU (Maas et al., 2013) and a BatchNorm Layer (Ioffe, 2015) to accelerate the model training. All methods are trained with Adagrad for optimization. We tuned the learning rate in 0.01, 0.005, 0.001. The batch size is set to 1024 and embedding dimension $d$ is set to 16.
Evaluate Metrics. we use two widely adopted metrics: AUC and GAUC. GAUC performs a weighted average over all users’ request group, and the weights are set to the number of samples of this user’s request..We will only apply GAUC on industrial dataset due to the number of groups is relatively small in public datasets.
Table 3. Hyper-parameters Analysis on Industrial Dataset
$\lambda=0.3$ $r=128$
LREA AUC GAUC LREA AUC GAUC
$r=32$ 0.7081 0.6222 $\lambda=0.05$ 0.7121 06298
$r=64$ 0.7097 0.6254 $\lambda=0.1$ 0.7138 06312
$r=128$ 0.7154 0.6343 $\lambda=0.3$ 0.7154 0.6343
- - $\lambda=0.5$ +0.7137 0.6315
- - $\lambda=0.8$ +0.7129 0.6297
4.2. Experiment Performance
Overall Perfomance. The evaluation results is shown in Table 2. SIM, ETA, SDIM, TWIN, outperform DIN. That indicates long-term behavior sequence is helpful for CTR prediction since it can cover more user’s historical interests. DIN(Long Sequence) has a stable improvement of AUC gain of 0.86%/0.65% compared with the best two-stage method on Amazon and Taobao dataset. We can also see a AUC gain of 0.49% and a GAUC gain of 0.79% in industrial dataset when comparing DIN(Long Sequence) with TWIN. But due to the limitation of inference latency, we can’t deploy DIN to long sequences. However, our LREA also outperforms all two-stage methods and it achieves similar results to long sequence target attention, and even exceeds DIN on industrial datasets but its inference complexity is much less than that of DIN(Long sequence).
Hyper-parameters Analysis. We conduct analysis experiments of rank $r$ and coefficient $\lambda$ on our industrial dataset. As shown in Table 3. First we fix $\lambda$ to 0.3, tune r from 32 to 128. The rank of $r=128$ achieves the best performance, and AUC/GAUC drop significantly when $r$ becomes smaller. This indicates that we can’t compression the representation of the original long sequences to a very low rank. We didn’t test $r$ 128, because although the larger r is, the better performance we get, there should be a appropriate rank to balance cache memory size and model performance. Next, we fix $r$ at 128 and adjust the value of $\lambda$ between 0.05 and 0.8. As mentioned earlier, $\lambda$ controls the nonlinearity and non-negativity in the compression module. If $\lambda$ is too large, it reduces the nonlinearity, resulting in a decline in model performance. Conversely, if $\lambda$ is too small, it skews the loss of the main task of CTR estimation. In our experiments, the model performs best when $\lambda=0.3$.
Online A/B Test. We conduct online A/B experiments to evaluate LREA in our online advertising system. The experiment lasts for 12 days and LREA achieves 5.88% CTR and 4.26% RPM gain compared to the product model (DIN+SIM). The inference rt only increases by 3 ms. And In comparison, the inference rt of DIN(Long sequence) will average increases about 30 ms, which causes 15% of user requests to time out.
5. Conclusion
In this paper, we introduced LREA, a novel attention mechanism that enhances CTR prediction in online advertising and recommendation systems. LREA effectively integrates long-term user behavior by using low-rank matrix decomposition and matrix absorption. This mitigates the performance degradation caused by traditional top K selection. Experimental results demonstrate that our model achieves significant improvements over the competitors.